# Automated Retinal Layer and Fluid Segmentation and Cross-sectional Analysis using Spectral Domain Optical Coherence Tomography Images for Diabetic Retinopathy

### v1

@@ Abstract
Background: Diabetic retinopathy (DR) is a leading microvascular complication of diabetes and a significant cause of vision loss. Without timely intervention, it can lead to irreversible blindness. Spectral Domain Optical Coherence Tomography (SD-OCT) enables high-resolution imaging of retinal structures, aiding in the early detection of pathological changes. AI-driven segmentation enhances diagnostic precision and reduces clinician workload. This study aims to develop a deep learning-based pipeline for automated retinal segmentation and thickness analysis to improve DR assessment and monitoring.
Methods: A deep neural network (DNN) was trained for automated segmentation of ten retinal layers, intra-retinal fluid, and hyperreflective foci (HRF). Initial manual annotations from five SD-OCT volumes were used to train the model, with iterative refinements improving accuracy. Performance was evaluated using five-fold cross-validation across four DNN architectures. Retinal thickness was measured via the K-nearest neighbours (KNN) algorithm, and Early Treatment Diabetic Retinopathy Study (ETDRS) maps visualized spatial variations. Statistical comparisons were made between Non-Proliferative (NPDR) and Proliferative Diabetic Retinopathy (PDR). The correlation between layer thickness and visual acuity within each DR group is also evaluated.
Results: SwinUNETR achieved the highest segmentation accuracy, with a Dice Similarity Coefficient (DSC) of 0.7719 and a Normalized Surface Dice (NSD) of 0.8149. VM-Unet showed superior DSC performance in specific layers (Vitreous, RNFL, GCL+IPL, ONL+IS), while SwinUNETR excelled in EZ, OS, RPE, and HRF regions. PDR patients exhibited increased thickness in the inner nasal and temporal of RNFL, whereas NPDR showed greater thickness in specific sectors of GCL+IPL, OS, and HRF sectors. Several sectors of INL and ONL+IS in NPDR show a significant correlation between layer thickening and better vision, but the reverse is true in PDR for RNFL, OPL, ONL+IS, and EZ. The accumulation of both fluid and HRF is also significantly correlated with vision decline in PDR, but no significance is found in NPDR. Similar findings can be derived from the predicted segmentation.
Conclusion: This study introduces an AI-based pipeline for automated DR assessment. It reduces manual annotation effort while maintaining high accuracy. The generated thickness maps provide valuable clinical insights, supporting disease monitoring and treatment planning.
@@ I Introduction
Diabetic retinopathy (DR) is a prevalent microvascular complication of diabetes mellitus (DM) and a leading cause of vision impairment worldwide (Klein, 2007). It is estimated that approximately 20% of diabetic individuals over the age of 50 will develop DR, which, if left untreated, can progress to severe visual impairment or blindness (Teo et al., 2021). The progression of DR is influenced by several risk factors, including prolonged diabetes duration, elevated glycated hemoglobin (HbA1c) levels, hypertension, hyperlipidemia, obesity, and smoking (Klein et al., 2014)(Zhou et al., 2025)(Mori et al., 2024)(Cheung et al., 2016)(Zhang et al., 2024)(Cao et al., 2017). Clinically, DR is categorized into two primary stages: Non-Proliferative Diabetic Retinopathy (NPDR) and Proliferative Diabetic Retinopathy (PDR). NPDR represents an early stage, often asymptomatic, characterized by microvascular abnormalities that progressively compromise retinal capillary integrity. Without timely medical intervention, NPDR can advance to PDR, a more severe stage marked by pathological neovascularization due to chronic retinal ischemia. This progression increases the risk of severe complications such as vitreous hemorrhage and retinal detachment, ultimately threatening vision.
Spectral Domain Optical Coherence Tomography (SD-OCT) is a cutting-edge, non-invasive imaging technique that provides high-resolution, cross-sectional visualization of retinal structures. Its real-time image acquisition capability makes it an invaluable tool for DR screening and early diagnosis (Gabriele E. Lang, 2007). SD-OCT enables the detection of subclinical retinal changes by quantifying variations in retinal thickness and identifying fluid accumulation. Accurate and reliable segmentation of retinal layers and pathological features is crucial for DR diagnosis and treatment planning. Figure 1 illustrates a representative SD-OCT B-scan with retinal layer and fluid segmentation. The segmentation delineates nine essential retinal layers: the Retinal Nerve Fiber Layer (RNFL), Ganglion Cell Layer and Inner Plexiform Layer (GCL+IPL), Inner Nuclear Layer (INL), Outer Plexiform Layer (OPL), Outer Nuclear Layer and Inner Segment Layer (ONL+IS), Ellipsoid Zone (EZ), Outer Segment Layer (OS), and Retinal Pigment Epithelium (RPE). The region above the Internal Limiting Membrane (ILM) is also identified as the Vitreous, while the Choroid lies beneath Bruch’s Membrane (BM). Fluid segmentation involves three primary fluid types: intraretinal fluid (IRF), subretinal fluid (SRF), and pigment epithelial detachment (PED), all of which appear as hypo-reflective regions between the ILM and BM. Furthermore, hyperreflective foci (HRF), which manifest as high-intensity dot-like or clustered lesions, are also segmented.
Figure 1: Example of retinal OCT segmentation(image on the right) for a single B-scan(image on the left). This image is acquired from a 59-year-old female patient with NPDR. Ten retinal layers are segmented from top to bottom, plus the fluid and HRF within the retinal body.
@@
Deep learning has emerged as a powerful alternative for automated retinal segmentation, offering greater robustness against variations in image quality and pathological abnormalities. Liu et al. utilized a ResNet-based convolutional neural network (CNN) combined with a random forest classifier for patch-wise layer segmentation (Liu et al., 2019). Kugelman et al. proposed a recurrent neural network (RNN) with a graph search framework to segment retinal layers in both healthy individuals and patients with age-related macular degeneration (AMD) (Kugelman et al., 2018). Hu et al. developed a multi-scale CNN capable of capturing different feature levels for improved segmentation accuracy (Hu et al., 2019). U-Net and its derivatives have become widely adopted among deep-learning models for medical image segmentation. U-Net’s encoder-decoder architecture, enhanced by skip connections, enables efficient spatial information preservation and mitigates vanishing gradient issues (Ronneberger et al., 2015). It has been successfully applied to retinal layer segmentation, fluid detection, and HRF analysis, achieving state-of-the-art (SOTA) performance (Roy et al., 2017)(Ma et al., 2021)(Tennakoon et al., 2018)(Schlegl et al., 2018). More recently, Vision Transformers (ViTs) have outperformed CNNs in large-scale datasets. Unlike CNNs, which rely on local receptive fields, ViTs employ self-attention mechanisms to capture global dependencies, which is particularly beneficial for detecting diffuse fluid regions. Xue et al. implemented a Swin-Transformer-based architecture for fluid segmentation in diabetic macular edema (DME) and AMD, demonstrating superior performance over traditional CNN-based models (Xue and Du, 2024). Kulyabin et al. leveraged the Segment Anything Model (SAM) for retinal fluid segmentation, incorporating point and bounding box prompts to outperform U-Net in macular hole and fluid segmentation tasks (Kulyabin et al., 2024). Despite these advancements, most existing studies focus on either the retinal layer or fluid segmentation, with limited efforts dedicated to integrating segmentation outcomes with clinical statistical analysis.
@@
Several studies have examined the relationship between retinal layer thickness, fluid accumulation, and DR severity. Browning et al. analyzed macular thickness across different DR severity levels and observed a correlation between macular thickening and increased risk of subclinical edema (Browning et al., 2008). Kim et al. investigated choroidal thickness alterations in DR and DME patients, reporting a significant increase in choroidal thickness as DR severity progressed from mild/moderate NPDR to PDR (Kim et al., 2013). Cho et al. assessed macular and peripapillary retinal thickness in DR subjects, identifying statistically significant differences in retinal thickness across seven anatomical regions between DR and control groups (Cho et al., 2010). Santos et al. demonstrated that fluid accumulation within the outer segment (OS) layer was significantly associated with central retinal thickness and visual impairment in DME patients (Santos T et al., 2024). These findings suggest that retinal layer thickness and fluid distribution are reliable biomarkers for DR diagnosis and progression monitoring.
@@
This study introduces a fully automated pipeline integrating retinal layer and fluid segmentation with a statistical analysis of structural changes in DR patients. The key contributions include:
@@
    Development of a scalable pipeline for large-scale manual segmentation refinement.
@@
    Quantitative evaluation of multiple SOTA deep learning models for retinal layer and fluid segmentation.
@@
    Cross-sectional analysis of retinal structural changes and demographic variations, providing clinically relevant insights into DR progression.
@@ II-A Data Acquisition
116 SD-OCT volumes were acquired from Sankara Nethralaya Eye Care Hospital in India. The imaging data was obtained using the Cirrus HD-OCT 5000 (Carl Zeiss Meditec, Dublin, CA, USA). Seventeen OCT volumes were captured in Macular Cube mode with a 512 × 128 pixels resolution, covering a 6 × 6 mm² macular region centered on the fovea. The remaining 99 volumes were scanned in OCTA mode at 350 × 350 pixels, covering a 6 × 6 mm² macular region. Despite differences in scanning speed and resolution, Wong et al. reported no significant variation in macular thickness measurements between the two modes (Wong et al., 2024). Table I presents the demographic details of the two DR groups, showing no significant differences in age, diabetes duration, or visual acuity (p $>$ 0.05). However, the gender distribution differs due to the limited number of female patients in the PDR group.
@@
TABLE I: Demographic information of experimental DR groups. Numerical values are presented as mean ± standard deviation (SD). The visual acuity is expressed in the logarithm of the Minimum Angle of Resolution (LogMAR). The p-values are calculated using the Student’s t-test.
@@ II-B Pre-processing
To prepare the raw OCT volumes for further analysis, we perform several pre-processing steps:
@@
  • The approximate retinal center in each B-scan is adjusted to align with the center along the axial direction. The axial retinal center is estimated by computing the average axial position of pixels whose axial intensity values are more significant than the lowest 20th percentile. This helps initialize a starting point for axial motion correction.
@@
  • A 3D Bounded Variation (BV) smoothing technique is applied to suppress noise while preserving sharp structural boundaries. This provides a better contrast between manual labelling and correction.
@@
  • Motion artifacts are corrected in each B-scan’s axial and lateral directions. Axial translations are determined through cross-registration using the moving average of the central B-scan as a reference. Lateral corrections are achieved by performing registration based on the adjacent B-scans’ discrete Fourier transform (DFT). Rotational adjustments are computed by transforming translations into polar space, using the moving average of the central B-scan as a reference.
II-C Manual Segmentation
@@
Figure 2 illustrates that the segmentation process follows a structured workflow. Initially, five volumes are manually annotated from scratch. To enhance efficiency, manual segmentation is performed on every fifth B-scan, while the intermediate B-scans are interpolated under the assumption that adjacent B-scans share structural similarities. However, B-scans that exhibit significant structural changes are individually labelled. These five manually labelled volumes serve as the initial training dataset for a deep neural network (DNN), which is trained using a data split ratio of 3:1:1 for training, validation, and testing. The trained model is then used to generate segmentation predictions for an additional 20 volumes, which are subsequently reviewed and manually corrected. This iterative process continues, with the network being retrained on an expanded dataset each time, ensuring that all volumes undergo accurate segmentation and manual verification.
@@
Figure 2: Manual segmentation pipeline. Multiple iterations were performed between DNN training and manual corrections. A thickness analysis was conducted after segmentation had been completed and verified.
II-D Networks
@@
We investigate the performance of four deep neural network (DNN) architectures, which are either widely used for medical image segmentation or have demonstrated state-of-the-art (SOTA) performance in related tasks:
@@
  • U-Net: A well-established medical image segmentation model that employs an encoder-decoder architecture with skip connections (Ronneberger et al., 2015). While effective, it may struggle with high-resolution inputs due to a lack of global contextual awareness.
@@
  • SegFormer: Designed by Xie et al., this framework utilizes a hierarchical transformer encoder combined with a lightweight MLP decoder to enhance feature extraction (Xie et al., 2021).
@@
  • SwinUNETR: Proposed by Hatamizadeh et al., this architecture replaces the CNN-based encoder in U-Net with a Swin Transformer encoder, enabling multi-scale feature extraction through a shifted windowing mechanism (Hatamizadeh et al., 2022).
@@
  • VM-UNet: Developed by Ruan et al., this model introduces a state-space model (SSM) and an asymmetric encoder-decoder structure, balancing computational efficiency while maintaining a global contextual view (Ruan et al., 2024).
The U-Net model is configured with a depth of five channels and three residual units. The 2D variant of SwinUNETR is used for training B-scans. The architectures of the remaining three networks are adopted as described in their original publications. The VM-UNet model is initialized using VMamba-S, pre-trained on the ImageNet-1k dataset (Liu et al., 2024).
II-E Training Settings

### v3

# Comprehensive Evaluation of OCT-based Automated Segmentation of Retinal Layer, Fluid and Hyper-Reflective Foci: Impact on Diabetic Retinopathy Severity Assessment

@@ Abstract
Background: Diabetic retinopathy (DR) is a major cause of vision loss, and early detection is essential to prevent irreversible blindness. Spectral Domain Optical Coherence Tomography (SD-OCT) enables high-resolution retinal imaging, while AI-driven segmentation improves diagnostic precision. However, segmentation performance varies across models, especially for DR cases with differing severity and complex fluid and hyperreflective foci (HRF) patterns. The clinical deployment of these models remains underexplored. This study develops an active-learning-based deep learning pipeline for automated segmentation of retinal layers, fluid, and HRF, comparing state-of-the-art (SOTA) models and evaluating their impact on DR assessment.
Methods: Four deep learning models (U-Net, SegFormer, SwinUNETR, VM-UNet) were trained on manually annotated SD-OCT volumes to segment ten retinal layers, fluid, and HRF. Five-fold cross-validation assessed segmentation performance. Retinal thickness was quantified using a K-nearest neighbours (KNN) algorithm and visualized via Early Treatment Diabetic Retinopathy Study (ETDRS) maps. Structural differences between Non-Proliferative (NPDR) and Proliferative DR (PDR) were analyzed, including correlations with visual acuity.
Results: SwinUNETR achieved the highest overall accuracy (DSC = 0.7719; NSD = 0.8149), while VM-UNet outperformed in specific layers. PDR showed increased OPL, fluid, and HRF thickness, whereas NPDR exhibited thickening in ONL+IS. In NPDR, thickening in RNFL, OPL, ONL+IS, and RPE correlated with reduced vision. In PDR, OS and EZ thickening and INL thinning were associated with visual impairment.
Conclusion: The proposed pipeline enables accurate, efficient DR analysis with reduced manual effort. SwinUNETR and VM-UNet performed robustly in complex regions, though HRF segmentation remains challenging. Thickness maps generated from auto-segmentation offer clinically relevant insights, supporting improved disease monitoring and treatment planning.
@@ I Introduction
Diabetic retinopathy (DR) is a prevalent microvascular complication of diabetes mellitus (DM) and a leading cause of vision impairment worldwide (Klein, 2007). It was estimated that approximately 20% of diabetic individuals over the age of 50 will develop DR, which, if left untreated, can progress to severe visual impairment or blindness (Teo et al., 2021). The progression of DR was influenced by several risk factors, including prolonged diabetes duration, elevated glycated hemoglobin (HbA1c) levels, hypertension, hyperlipidemia, obesity, and smoking (Klein et al., 2014)(Zhou et al., 2025)(Mori et al., 2024)(Cheung et al., 2016)(Zhang et al., 2024)(Cao et al., 2017). Clinically, DR is categorized into two primary stages: Non-Proliferative Diabetic Retinopathy (NPDR) and Proliferative Diabetic Retinopathy (PDR). NPDR represents an early stage, often asymptomatic, characterized by microvascular abnormalities that progressively compromise retinal capillary integrity. Without timely medical intervention, NPDR can advance to PDR, a more severe stage marked by pathological neovascularization due to chronic retinal ischemia. This progression increases the risk of severe complications such as vitreous hemorrhage and retinal detachment, ultimately threatening vision.
Spectral Domain Optical Coherence Tomography (SD-OCT) is a cutting-edge, non-invasive imaging technique that provides high-resolution, cross-sectional visualization of retinal structures. Its real-time image acquisition capability makes it an invaluable tool for DR screening and early diagnosis (Gabriele E. Lang, 2007). SD-OCT enables the detection of subclinical retinal changes by quantifying variations in retinal thickness and identifying fluid accumulation. Accurate and reliable segmentation of retinal layers and pathological features is crucial for DR diagnosis and treatment planning.
@@
Deep learning has emerged as a powerful alternative for automated retinal segmentation, offering greater robustness against variations in image quality and pathological abnormalities. Liu et al. utilized a ResNet-based convolutional neural network (CNN) combined with a random forest classifier for patch-wise layer segmentation (Liu et al., 2019). Kugelman et al. proposed a recurrent neural network (RNN) with a graph search framework to segment retinal layers in both healthy individuals and patients with age-related macular degeneration (AMD) (Kugelman et al., 2018). Hu et al. developed a multi-scale CNN capable of capturing different feature levels for improved segmentation accuracy (Hu et al., 2019). U-Net and its derivatives have become widely adopted among deep-learning models for medical image segmentation. U-Net’s encoder-decoder architecture, enhanced by skip connections, enables efficient spatial information preservation and mitigates vanishing gradient issues (Ronneberger et al., 2015). It has been successfully applied to retinal layer segmentation, fluid detection, and HRF analysis, achieving state-of-the-art (SOTA) performance (Roy et al., 2017, Ma et al., 2021, Tennakoon et al., 2018, Schlegl et al., 2018). Generative adversarial networks (GANs) were also used for retinal boundary augmentation and segmentation adaptation cross multiple OCT domains(Kugelman et al., 2023, Chen et al., 2023). Vision Transformers (ViTs) have recently outperformed CNNs in large-scale datasets. Unlike CNNs, which rely on local receptive fields, ViTs employ self-attention mechanisms to capture global dependencies, which is particularly beneficial for detecting diffuse fluid regions. Xue et al. implemented a Swin-Transformer-based architecture for fluid segmentation in diabetic macular edema (DME) and AMD, demonstrating superior performance over traditional CNN-based models (Xue and Du, 2024). Kulyabin et al. leveraged the Segment Anything Model (SAM) for retinal fluid segmentation, incorporating point and bounding box prompts to outperform U-Net in macular hole and fluid segmentation tasks (Kulyabin et al., 2024). Despite these advancements, most existing studies focus on either the retinal layer or fluid segmentation, with varying levels of segmentation performance on pathological clinical features. However, limited efforts are dedicated to investigating the effect of automated segmentation performance on NPDR/PDR classification or prognosis, which is crucial to evaluating their clinical translation.
@@
Studies have examined the relationship between retinal layer thickness, fluid accumulation, and DR severity. Browning et al. analyzed macular thickness across different DR severity levels and observed a correlation between macular thickening and increased risk of subclinical edema (Browning et al., 2008). Kim et al. investigated choroidal thickness alterations in DR and DME patients, reporting a significant increase in choroidal thickness as DR severity progressed from mild/moderate NPDR to PDR (Kim et al., 2013). Cho et al. assessed macular and peripapillary retinal thickness in DR subjects, identifying statistically significant differences in retinal thickness across seven anatomical regions between DR and control groups (Cho et al., 2010). Santos et al. demonstrated that fluid accumulation within the outer segment (OS) layer is significantly associated with central retinal thickness and visual impairment in DME patients (Santos T et al., 2024). These findings suggest that retinal layer thickness and fluid distribution are both reliable biomarkers for DR diagnosis and progression monitoring. However, limited efforts are dedicated to investigating the effect of automated segmentation performance on DR/PDR classification or prognosis, which is crucial to evaluating their clinical translation.
@@
The current study introduces an end-to-end framework integrating retinal layer and fluid segmentation with a statistical analysis of structural changes in DR patients. The key contributions include:
@@
    Development of an efficient active-learning-based segmentation pipeline for severely pathological DR patients.
@@
    Comprehensive evaluation of multiple SOTA deep learning models, revealing differential performance on segmenting retinal layers, fluid, and HRF segmentations, using both volume- and thickness-based evaluation metrics for all biomarkers, differentially considering under-segmentation and over-segmentation cases.
@@
    Evaluate the clinical translatability of the auto-segmentation-based retinal thicknesses, fluid and HRF biomarkers for differentiating DR severity, as well as their association with visual acuity.
@@ II-A Data Acquisition
116 SD-OCT volumes were acquired from Sankara Nethralaya Eye Care Hospital in India. The imaging data was obtained using the Cirrus HD-OCT 5000 (Carl Zeiss Meditec, Dublin, CA, USA). Seventeen OCT volumes were captured in Macular Cube mode with a 512 × 128 pixels resolution, with the remaining 99 volumes scanned in OCTA mode at 350 × 350 pixels. Both modes covered a 6 × 6 mm² macular region centered on the fovea. Despite differences in scanning speed and resolution, Wong et al. reported no significant variation in macular thickness measurements between the two modes (Wong et al., 2024). Table I presents the demographic details of the subjects in two DR severity groups, showing no significant differences in age, diabetes duration, or visual acuity (p $>$ 0.05). However, the gender distribution differs due to the limited number of female patients in the PDR group. The variance inflation factor (VIF) is calculated for DR groups, age, gender, duration of diabetes, and visual acuity. No significant multicollinearity is found as all values are close to 1.
@@
VIF - 1.158384 1.840780 1.162841 1.642617
TABLE I: Demographic information of experimental DR groups. Numerical values are presented as mean ± standard deviation (SD). The visual acuity is expressed in the logarithm of the Minimum Angle of Resolution (LogMAR). The p-values are calculated using the Welch’s t-test. The variance inflation factor (VIF) is calculated for each variable including the DR category.
@@ II-B Pre-processing
To prepare the raw OCT volumes for further analysis, we performed several pre-processing steps:
@@
  • The approximate retinal center in each B-scan was adjusted to align with the center along the axial direction. The axial retinal center was estimated by computing the average axial position of pixels whose axial intensity values are more significant than the lowest 20th percentile. This helped initialize a starting point for axial motion correction.
@@
  • A 3D Bounded Variation (BV) smoothing technique was applied to suppress noise while preserving smoother structural boundaries, providing better contrasts for manual labelling and model prediction.
@@
  • Motion artifacts among adjacent B-scan were corrected in both the axial and lateral directions. Axial translations were determined through cross-registration using the moving average of the central B-scan as a reference. Lateral corrections were achieved by performing registration based on the adjacent B-scans’ discrete Fourier transform (DFT). Rotational adjustments were computed by transforming translations into polar space, using the moving average of the central B-scan as a reference.
II-C Active-Learning-Based Ground Truth Segmentation Annotation
@@
Figure 1 illustrates that the active-learning-based semi-automatic segmentation follows a structured human-in-the-loop (HITL) interactive labelling workflow. SwinUNETR was used as the backbone network architecture for the active-learning workflow. The choice of architecture will only affect the efficiency of the manual labelling process, but it will not result in discrepancies in quantitative evaluations. Initially, five volumes were manually annotated from scratch. Manual segmentation was performed on every fifth B-scan, while the intermediate B-scans were interpolated under the assumption that adjacent B-scans share structural similarities. However, B-scans that exhibit significant structural changes were individually labelled and corrected. These five manually labelled volumes served as the first iteration of the training dataset to train a deep neural network (DNN) with a data split ratio of 3:1:1 for training, validation, and testing. The initially trained model was then used to generate segmentation predictions for an additional 20 volumes, which were subsequently reviewed and manually corrected. This iterative process continued with a 7:2:1 data split ratio in the subsequent iterations for training, validation, and testing, with the network being retrained on an expanded dataset each time, ensuring that all volumes undergo accurate segmentation and manual verification. Volume splits were stratified to ensure pathological cases with all label types (i.e. retinal layer, fluid, and HRF) presented in training, validation, and testing sets,
@@
Figure 1: Manual segmentation pipeline. Multiple iterations were performed between DNN training and manual corrections. A thickness analysis was conducted after segmentation has been completed and verified.
II-D Automatic Segmentation Networks Architecture
@@
We investigated the performance of four deep neural network (DNN) architectures, which are either widely used for medical image segmentation or have demonstrated SOTA performance in related tasks:
@@
  • U-Net: The most widely-used well-established medical image segmentation model that employs a CNN-based encoder-decoder architecture with skip connections (Ronneberger et al., 2015). While effective, it may struggle with high-resolution inputs due to a lack of global contextual awareness. The U-Net model is configured with a depth of five channels and three residual units.
@@
  • SegFormer: A transformer-based architecture proposed by Xie et al., which utilizes a hierarchical transformer encoder combined with a lightweight MLP decoder to enhance feature extraction (Xie et al., 2021). The 2D variant of SwinUNETR is used for training B-scans.
@@
  • SwinUNETR: A CNN-Transformer-composited architecture proposed by Hatamizadeh et al., which replaces the CNN-based encoder in U-Net with a Swin Transformer encoder, enabling multi-scale feature extraction through a shifted windowing mechanism. (Hatamizadeh et al., 2022). This is also the architecture that is used for the semi-automatic generation of the ground truth segmentation labels through the HITL active-learning process.
@@
  • VM-UNet: A recently proposed novel architecture developed by Ruan et al., this model introduces a state-space model (SSM) and an asymmetric encoder-decoder structure. It models the visual data as an evolving state, efficiently capturing both local and global dynamics with a structure inspired by continuous dynamic systems, balancing computational efficiency while maintaining a global contextual view (Ruan et al., 2024).
II-E Segmentation Model Training